1) Al + Pb(NO3)2 --->
2) Cl2 + NaI --->
3) Fe + AgC2H3O2 --->
4) Al + CuCl2 --->
5) Br2 + CaI2 --->
6) Al + HCl --->
7) Mg + HCl --->
8) Zn + H2SO4 --->
9) Fe + CuSO4 --->
10) Cl2 + MgI2 --->
1
Expert's answer
2020-04-03T11:51:59-0400

1)2Al + 3Pb(NO3)2 = 2Al(NO3)3 + 3Pb

2)Cl2 + 2NaI = 2NaCl + I2

3)Fe + 2AgC2H3O2 = Fe(C2H3O2)2 + 2Ag

4)2Al + 3CuCl2 = 2AlCl3 + 3Cu

5)Br2 + CaI2 = CaBr2 + I2

6)2Al + 6HCl = 2AlCl3 + 3H2

7)Mg + 2HCl = MgCl2 + H2

8)Zn + H2SO4 = ZnSO4 + H2

9)Fe + CuSO4 = FeSO4 + Cu

10)Cl2 + MgI2 = MgCl2 + I2